# RawHash2: Mapping Raw Nanopore Signals Using Hash-Based Seeding and Adaptive Quantization

### v4

@@ Abstract
Summary: Raw nanopore signals can be analyzed while they are being generated, a process known as real-time analysis. Real-time analysis of raw signals is essential to utilize the unique features that nanopore sequencing provides, enabling the early stopping of the sequencing of a read or the entire sequencing run based on the analysis. The state-of-the-art mechanism, RawHash, offers the first hash-based efficient and accurate similarity identification between raw signals and a reference genome by quickly matching their hash values. In this work, we introduce RawHash2, which provides major improvements over RawHash, including a more sensitive quantization and chaining algorithms, weighted mapping decisions, frequency filters to reduce ambiguous seed hits, minimizers for hash-based sketching, and support for the R10.4 flow cell version and POD5 and SLOW5 file formats. Compared to RawHash, RawHash2 provides better F1 accuracy (on average by $10.57\%$ and up to $20.25\%$) and better throughput (on average by 4.0×4.0$\times$4.0 × and up to $\times9$) than RawHash.
@@ 1 Introduction
Computational techniques that can analyze the raw signals while they are generated at a speed that matches the throughput of nanopore sequencing are called real-time analysis. Figure 1 shows the two unique benefits that real-time analysis offers. First, real-time analysis allows for overlapping sequencing time with analysis time as raw signals can be analyzed while they are being generated. Second, computational mechanisms can stop the sequencing of a read or the entire sequencing run early without sequencing the entire molecule or the sample using techniques known as Read Until [6] and Run Until [7]. The development of accurate and fast mechanisms for real-time analysis has the potential to significantly reduce the time and cost of genome analysis.
@@
Among the existing works, RawHash [5] is the state-of-the-art mechanism that can accurately perform real-time mapping of raw nanopore signals for large genomes without translating them to bases with a hash-based seed-and-extend mechanism [18]. Despite its strengths in accuracy and performance, particularly for large genomes like the human genome, RawHash exhibits several limitations that require further improvements. First, RawHash utilizes a simple quantization algorithm that assumes the raw signals are distributed uniformly across their normalized value range, which limits its efficiency and accuracy. Second, RawHash uses chaining algorithm similar to that used in Sigmap [3] without incorporating penalty scores used in minimap2 [19], which constrains its ability for more sensitive mapping. Third, RawHash performs chaining on all seed hits without filtering any of these seed hits, which substantially increases the workload of the chaining algorithm due to a large number of seed hits to chain. Fourth, the decision-making mechanism in RawHash for mapping reads to a reference genome in real-time relies on one of the mapping conditions being true (e.g., the ratio between the best and second-best chain scores), which makes it more prone to the outliers that can satisfy one of these conditions. A more robust and statistical approach that incorporates features beyond chaining scores can provide additional insights for making more sensitive and quick mapping decisions. Fifth, while the hash-based mechanism in RawHash is compatible with existing sketching techniques such as minimizers [20, 19], strobemers [21], and fuzzy seed matching as in BLEND [22], the benefits of these techniques are unknown for raw signal analysis as they are not used in RawHash. Such evaluations could potentially provide additional insights on how to use the existing hash-based sketching techniques and reduce storage requirements while maintaining high accuracy. Sixth, RawHash lacks the support for recent advancements such as the newer R10.4 flow cell version. The integration of these features can accelerate the adoption for both real-time and offline analysis.
@@ 2 Methods
RawHash is a mechanism to perform mapping between raw signals by quickly matching their hash values. We provide the details of the RawHash mechanism in Supplementary Section S1. RawHash2 provides substantial improvements over RawHash in six key directions. First, to generate more accurate and distinct hash values from raw signals, RawHash2 improves the quantization mechanism with an adaptive approach such that signal values are quantized non-uniformly based on the characteristics of a nanopore model. Second, to provide more accurate mapping, RawHash2 improves the chaining algorithm in RawHash with more accurate penalty scores. Third, to reduce the workload in chaining for improved performance, we integrate a frequency filter to quickly eliminate the seed hits that occur too frequently. Fourth, to make more accurate and quick mapping decisions, RawHash2 determines whether a read should be mapped at a specific point during sequencing by using a weighted sum of multiple features. Fifth, to reduce the storage requirements of seeds, RawHash2 incorporates and evaluates the benefits of minimizer sketching technique. Sixth, RawHash2 includes support for the latest features introduced by ONT, such as new file formats and flow cells.
@@ 3.1 Evaluation Methodology
To estimate the benefits in sequencing time and the cost per read, we identify the average sequencing length before making the mapping decision for a read. For all of our analyses, we use the default parameters of each tool as we show in Supplementary Table S7. Supplementary Table S6 shows the real dataset details we use in our evaluation, including more details about sequencing run settings and flow cell versions (i.e., R9.4.1 and R10.4). For all the datasets except D7, we use already basecalled reads available with the raw electrical signals. For the D7 dataset, we basecall the raw signals using the Dorado basecaller. Although RawHash2 does not use the minimizer sketching technique by default to achieve the maximum accuracy, we evaluate the benefits of minimizers in RawHash2, which we refer to as RawHash2-Minimizer. Since RawHash2 is the only tool that can perform real-time mapping of raw R10.4 nanopore signals, we show the corresponding results when using the R10.4 dataset without comparing these results with the existing tools. When comparing RawHash2 to other tools we always use FAST5 files containing raw signals from R9.4 flow cells on an isolated machine and SSD. We use AMD EPYC 7742 processor at 2.26GHz to run the tools. We use 32 threads for all the tools.
@@ 3.2 Throughput
Figure 2 shows the results for 1) throughput per single CPU thread and 2) number of pores that a single CPU thread can analyze as annotated by the values inside the bars. We make three key observations. First, we find that RawHash2 provides average throughput 26.5×26.5$\times$26.5 ×, 19.2×19.2$\times$19.2 ×, and 4.0×4.0$\times$4.0 × better than UNCALLED, Sigmap, and RawHash, respectively. Such a speedup, specifically over the earlier work RawHash, is achieved by reducing the workload of chaining with the unique and accurate hash values using the new quantization mechanism and the filtering technique see the filtering ratios in Supplementary Table S5). Second, we find that RawHash2-Minimizer enables reducing the computational requirements for mapping raw signals and enables improving the average throughput by 2.5×2.5$\times$2.5 × compared to RawHash2, while the other computational resources, such as the peak memory usage and CPU time in both indexing and mapping, and the mean time spent per read are also significantly reduced as shown in Supplementary Tables S3 and Supplementary Figure S3. Third, RawHash2-Minimizer requires at most 7 threads for analyzing the entire flowcell for any evaluated dataset, while RawHash2 requires at most 2 threads for smaller genomes and 9 to 26 threads for Green Algae and human. This shows that RawHash2 and RawHash2-Minimizer can reduce computational requirements and energy consumption significantly compared to 28 threads required, on average, regardless of the genome size for UNCALLED, which is critical for portable sequencing. We conclude that RawHash2 and RawHash2-Minimizer significantly reduce the computational overhead of mapping raw signals to reference genomes, enabling better scalability to even larger genomes.
@@ 3.4 Sequencing Time and Cost
In Supplementary Table S4 and S2, we show the results when using different file formats for storing raw signals (i.e., FAST5, POD5, and BLOW5) and R10.4, respectively. We make two key observations. First, we find that POD5 significantly speeds up total elapsed time, especially for multi-threaded analysis. Thread utilization increases significantly with faster read and write operations, shifting the bottleneck in the entire workload from the read and write operations to multi-threaded mapping computations. Second, we find that RawHash2 can perform accurate and fast analysis when using raw signals from R10.4, although RawHash2 achieves lower accuracy with R10.4 than using R9.4. This is likely because 1) we use a k-mer model optimized for the R10.4.1 flow cell version rather than R10.4, and 2) minimap2 can provide more accurate mapping due to improved accuracy of these basecalled reads. Future work can focus on generating a k-mer model specifically designed for R10.4 to generate more accurate results. We exclude the accuracy results for R10.4.1 as the number of events found for R10.4.1 is around $35\%$ larger than that of R10.4, which leads to inaccurate mapping. We suspect that our segmentation algorithm and parameters are not optimized for R10.4.1. Our future work will focus on improving these segmentation parameters and techniques to achieve higher accuracy with R10.4.1 as well as RNA sequencing data. We believe this can be achieved because RawHash2 1) is highly flexible to change all the parameters corresponding to segmentation and 2) can map accurately without requiring long sequencing lengths (Table 2), which can mainly be useful for RNA read sets.
We conclude that RawHash2 is the only work that can provide accurate and fast analysis when using the recent features released by ONT.
@@ 4 Conclusion
S2 Accuracy
S2.1 Read Mapping Accuracy
In Supplementary Table S1, we show the read mapping accuracy in all metrics (i.e., F1, Precision, and Recall) for all datasets. In Figure S2, we show the same results as reported in Supplementary Table S1 for visualizing the comparisons between tools and the trade-offs between each accuracy metrics in all datasets.
@@
S3 Performance
S3.1 Runtime, Peak Memory Usage, and Throughput
@@
S3.2 Impact of Different File Formats on Performance
@@
Table S4: Comparison of run time and thread utilization when using different file formats in RawHash2.
@@
Real Time (sec)
RawHash2-FAST5 279 80
RawHash2-POD5 99 18
RawHash2-BLOW5 290 42
Thread Utilization (%, max. 32 threads)
RawHash2-FAST5 4.63 6.77
RawHash2-POD5 11.51 27.4
RawHash2-BLOW5 5.42 14.37
S3.3 Mapping Time per Read
@@
S3.4 Combined benefits of performance, accuracy, and average sequencing length
@@
S3.5 Ratio of Filtered Seeds from Frequency Filter
Supplementary Table S5 shows the ratio of seed matches filtered out by the frequency filtered in RawHash2.
@@
S4 Configuration
S4.1 Datasets
In Supplementary Table S6 we show the details of the datasets used in our evaluation along with their corresponding sequencing run settings.
@@
D1 SARS-CoV-2 MinION R9.4.1 e8 (FLO-MIN106) 450 4000 Guppy HAC 1,382,016 594M CADDE Centre GCF_009858895.2 29,903
D2 E. coli GridION R9.4.1 e8 (FLO-MIN106) 450 4000 Guppy HAC 353,317 2,365M ERR9127551 GCA_000007445.1 5M
@@
D5 Human MinION R9.4.1 e8 (FLO-MIN106) 450 4000 Guppy Flip-Flop 269,507 1,584M FAB42260 T2T-CHM13 (v2) 3,117M
D6 E. coli GridION R10.4 e8.1 (FLO-MIN112) 450 4000 Guppy HAC 1,172,775 6,123M ERR9127552 GCA_000007445.1 5M
D7 S. aureus GridION R10.4 e8.1 (FLO-MIN112) 450 4000 Dorado SUP 407,727 1,281M SRR21386013 GCF_000144955.2 2.8M
@@
S4.3 Versions
Supplementary Table S9 shows the version and the link to these corresponding versions of each tool and library that we use in our experiments and in RawHash2, respectively.

### v5

@@ Abstract
Summary: Raw nanopore signals can be analyzed while they are being generated, a process known as real-time analysis. Real-time analysis of raw signals is essential to utilize the unique features that nanopore sequencing provides, enabling the early stopping of the sequencing of a read or the entire sequencing run based on the analysis. The state-of-the-art mechanism, RawHash, offers the first hash-based efficient and accurate similarity identification between raw signals and a reference genome by quickly matching their hash values. In this work, we introduce RawHash2, which provides major improvements over RawHash, including more sensitive quantization and chaining algorithms, weighted mapping decisions, frequency filters to reduce ambiguous seed hits, minimizers for hash-based sketching, and support for the R10.4 flow cell version and POD5 and SLOW5 file formats. Compared to RawHash, RawHash2 provides better F1 accuracy (on average by $10.57\%$ and up to $20.25\%$) and better throughput (on average by 4.0×4.0$\times$4.0 × and up to $\times9$) than RawHash.
@@ 1 Introduction
Computational techniques that can analyze the raw signals while they are generated at a speed that matches the throughput of nanopore sequencing are called real-time analysis. Figure 1 shows the two unique benefits that real-time analysis offers. First, real-time analysis allows for overlapping sequencing time with analysis time, as raw signals can be analyzed while they are being generated. Second, computational mechanisms can stop the sequencing of a read or the entire sequencing run early without sequencing the entire molecule or the sample using techniques known as Read Until [6] and Run Until [7]. The development of accurate and fast mechanisms for real-time analysis has the potential to significantly reduce the time and cost of genome analysis.
@@
Among the existing works, RawHash [5] is the state-of-the-art mechanism that can accurately perform real-time mapping of raw nanopore signals for large genomes without translating them to bases with a hash-based seed-and-extend mechanism [18]. Despite its strengths in accuracy and performance, particularly for large genomes like the human genome, RawHash exhibits several limitations that require further improvements. First, RawHash utilizes a simple quantization algorithm that assumes the raw signals are distributed uniformly across their normalized value range, which limits its efficiency and accuracy. Second, RawHash uses a chaining algorithm similar to that used in Sigmap [3] without incorporating penalty scores used in minimap2 [19], which constrains its ability for more sensitive mapping. Third, RawHash performs chaining on all seed hits without filtering any of these seed hits, which substantially increases the workload of the chaining algorithm due to a large number of seed hits to chain. Fourth, the decision-making mechanism in RawHash for mapping reads to a reference genome in real-time relies on one of the mapping conditions being true (e.g., the ratio between the best and second-best chain scores), which makes it more prone to the outliers that can satisfy one of these conditions. A more robust and statistical approach that incorporates features beyond chaining scores can provide additional insights for making more sensitive and quick mapping decisions. Fifth, while the hash-based mechanism in RawHash is compatible with existing sketching techniques such as minimizers [20, 19], strobemers [21], and fuzzy seed matching as in BLEND [22], the benefits of these techniques are unknown for raw signal analysis as they are not used in RawHash. Such evaluations could potentially provide additional insights on how to use the existing hash-based sketching techniques and reduce storage requirements while maintaining high accuracy. Sixth, RawHash lacks the support for recent advancements, such as the newer R10.4 flow cell version. The integration of these features can accelerate the adoption of both real-time and offline analysis.
@@ 2 Methods
RawHash is a mechanism to perform mapping between raw signals by quickly matching their hash values. We provide the details of the RawHash mechanism in Supplementary Section A. RawHash2 provides substantial improvements over RawHash in six key directions. First, to generate more accurate and distinct hash values from raw signals, RawHash2 improves the quantization mechanism with an adaptive approach such that signal values are quantized non-uniformly based on the characteristics of a nanopore model. Second, to provide more accurate mapping, RawHash2 improves the chaining algorithm in RawHash with more accurate penalty scores. Third, to reduce the workload in chaining for improved performance, we integrate a frequency filter to quickly eliminate the seed hits that occur too frequently. Fourth, to make more accurate and quick mapping decisions, RawHash2 determines whether a read should be mapped at a specific point during sequencing by using a weighted sum of multiple features. Fifth, to reduce the storage requirements of seeds, RawHash2 incorporates and evaluates the benefits of minimizer sketching technique. Sixth, RawHash2 includes support for the latest features introduced by ONT, such as new file formats and flow cells.
@@ 3.1 Evaluation Methodology
To estimate the benefits in sequencing time and the cost per read, we identify the average sequencing length before making the mapping decision for a read. For all of our analyses, we use the default parameters of each tool as we show in Supplementary Table S7. Supplementary Table S6 shows the real dataset details we use in our evaluation, including more details about sequencing run settings and flow cell versions (i.e., R9.4.1 and R10.4). For all the datasets except D7, we use already basecalled reads available with the raw electrical signals. For the D7 dataset, we basecall the raw signals using the Dorado basecaller. Although RawHash2 does not use the minimizer sketching technique by default to achieve the maximum accuracy, we evaluate the benefits of minimizers in RawHash2, which we refer to as RawHash2-Minimizer. Since the evaluated versions of UNCALLED, Sigmap, and RawHash do not provide the support for R10.4 dataset, we show the corresponding results with the R10.4 dataset without comparing to these tools. When comparing RawHash2 to other tools we always use FAST5 files containing raw signals from R9.4 flow cells on an isolated machine and SSD. We use AMD EPYC 7742 processor at 2.26GHz to run the tools. We use 32 threads for all the tools.
@@ 3.2 Throughput
Figure 2 shows the results for 1) throughput per single CPU thread and 2) number of pores that a single CPU thread can analyze as annotated by the values inside the bars. We make three key observations. First, we find that RawHash2 provides average throughput 26.5×26.5$\times$26.5 ×, 19.2×19.2$\times$19.2 ×, and 4.0×4.0$\times$4.0 × better than UNCALLED, Sigmap, and RawHash, respectively. Such a speedup, specifically over the earlier work RawHash, is achieved by reducing the workload of chaining with the unique and accurate hash values using the new quantization mechanism and the filtering technique (see the filtering ratios in Supplementary Table S5). Second, we find that RawHash2-Minimizer enables reducing the computational requirements for mapping raw signals and enables improving the average throughput by 2.5×2.5$\times$2.5 × compared to RawHash2, while the other computational resources, such as the peak memory usage and CPU time in both indexing and mapping, and the mean time spent per read are also significantly reduced as shown in Supplementary Tables S3 and Supplementary Figure S3. Third, RawHash2-Minimizer requires at most 7 threads for analyzing the entire flowcell for any evaluated dataset, while RawHash2 requires at most 2 threads for smaller genomes and 9 to 26 threads for Green Algae and human. This shows that RawHash2 and RawHash2-Minimizer can reduce computational requirements and energy consumption significantly compared to 28 threads required, on average, regardless of the genome size for UNCALLED, which is critical for portable sequencing. We conclude that RawHash2 and RawHash2-Minimizer significantly reduce the computational overhead of mapping raw signals to reference genomes, enabling better scalability to even larger genomes.
@@ 3.4 Sequencing Time and Cost
In Supplementary Table S4 and S2, we show the results when using different file formats for storing raw signals (i.e., FAST5, POD5, and BLOW5) and R10.4, respectively. We make two key observations. First, we find that POD5 and SLOW5 significantly speed up total elapsed time compared to FAST5. These results indicate that a large portion of the overhead spent for reading from a file can be mitigated with approaches that can perform faster compression and decompression, as these signal files are mostly stored in a compressed form. Second, we find that RawHash2 can perform fast analysis with reasonable accuracy that can be useful for certain use cases (e.g., contamination analysis) when using raw signals from R10.4, although RawHash2 achieves lower accuracy with R10.4 than using R9.4. This is likely because 1) we use a k-mer model optimized for the R10.4.1 flow cell version rather than R10.4, and 2) minimap2 can provide more accurate mapping due to improved accuracy of these basecalled reads. Future work can focus on generating a k-mer model specifically designed for R10.4 to generate more accurate results. We exclude the accuracy results for R10.4.1 as the number of events found for R10.4.1 is around $35\%$ larger than that of R10.4, which leads to inaccurate mapping. We suspect that our segmentation algorithm and parameters are not optimized for R10.4.1. Our future work will focus on improving these segmentation parameters and techniques to achieve higher accuracy with R10.4.1 as well as RNA sequencing data. We believe this can be achieved because RawHash2 1) is highly flexible to change all the parameters corresponding to segmentation and 2) can map accurately without requiring long sequencing lengths (Table 2), which can mainly be useful for RNA read sets. We conclude that RawHash2 can provide accurate and fast analysis when using the recent features released by ONT.
@@ 4 Conclusion
Acknowledgments
We thank all members of the SAFARI Research Group for the stimulating and scholarly intellectual environment they provide. We acknowledge the generous gift funding provided by our industrial partners (especially by Google, Huawei, Intel, Microsoft, VMware), which has been instrumental in enabling the decade+ long research we have been conducting on accelerating genome analysis. This work is also partially supported by the Semiconductor Research Corporation (SRC), the European Union’s Horizon programme for research and innovation [101047160 - BioPIM] and the Swiss National Science Foundation (SNSF) [200021_213084].
@@ A RawHash Overview
B Accuracy
B.1 Read Mapping Accuracy
In Supplementary Table S1, we show the read mapping accuracy in all metrics (i.e., F1, Precision, and Recall) for all datasets. In Figure S2, we show the same results as reported in Supplementary Table S1 for visualizing the comparisons between tools and the trade-offs between each accuracy metric in all datasets.
@@ B.1 Read Mapping Accuracy
C Performance
C.1 Runtime, Peak Memory Usage, and Throughput
@@
C.2 Impact of Different File Formats on Performance
@@
We note that even if we use multiple threads when running RawHash2, the file IO step (i.e., reading from or writing to a file) always uses a single thread and is overlapped with the mapping step (i.e., either the read or write operation is run in parallel together with the mapping step by using one thread where the mapping step takes rest of the allocated threads). The design is due to the pipelining implementation strategy we adopt, similar to the minimap2 implementation [19]. We note that if RawHash2 is run using a single thread, none of these steps overlap with each other, and they run sequentially using only one thread, which is our evaluation setting we show in Supplementary Table S4.
Table S4: Comparison of overall execution time when using different file formats in RawHash2 in a single-threaded mode.
@@
Elapsed Time (mm:ss)
RH2-FAST5 19:27 08:35
RH2-POD5 16:55 07:33
RH2-BLOW5 17:32 07:38
RH2-Min.-FAST5 12:13 03:56
RH2-Min.-POD5 09:42 02:56
RH2-Min.-BLOW5 10:16 03:02
C.3 Mapping Time per Read
@@
C.4 Combined benefits of performance, accuracy, and average sequencing length
@@
C.5 Ratio of Filtered Seeds from Frequency Filter
Supplementary Table S5 shows the ratio of seed hits filtered out by the frequency filtered in RawHash2. We calculate these ratios in three steps. First, for each seed (i.e., a hash value that RawHash2 constructs from raw signals), we perform a query to the hash table that is used as an index. If the hash value exists, the table returns a list of genomic regions that share the same hash value. Each region counts as a seed hit, and the list length indicates the number of seed hits. Second, for all seeds generated from raw signals, we count 1) the overall number of seed hits and 2) the number of seed hits filtered out by frequency filter. We note that if the list length (i.e., number of seed hits) returned after querying a particular seed is above a certain threshold (defined by our frequency filter), all seed hits within the same list are filtered out. Third, we calculate the ratio of filtered seed hits to the total seed hits and report these ratios in Supplementary Table S5.
@@ C.5 Ratio of Filtered Seeds from Frequency Filter
D Configuration
D.1 Datasets
In Supplementary Table S6 we show the details of the datasets used in our evaluation and their corresponding sequencing run settings. The Basecaller Model column shows the details about the basecaller model and the version we use. Except for the D7 dataset, all other datasets include the basecalled sequences within their corresponding FAST5 files or the corresponding accession numbers available at NCBI. We provide the scripts to extract these basecalled sequences on the GitHub page of RawHash2. For the D7 dataset, we provide the necessary commands to run dorado for basecalling on the GitHub page.
@@ D.1 Datasets
D1 SARS-CoV-2 MinION R9.4.1 e8 (FLO-MIN106) 450 4000 Guppy HAC v3.2.6 1,382,016 594M CADDE Centre GCF_009858895.2 29,903
D2 E. coli GridION R9.4.1 e8 (FLO-MIN106) 450 4000 Guppy HAC v5.0.12 353,317 2,365M ERR9127551 GCA_000007445.1 5M
@@
D5 Human MinION R9.4.1 e8 (FLO-MIN106) 450 4000 Guppy Flip-Flop v2.3.8 269,507 1,584M FAB42260 T2T-CHM13 (v2) 3,117M
D6 E. coli GridION R10.4 e8.1 (FLO-MIN112) 450 4000 Guppy HAC v5.0.16 1,172,775 6,123M ERR9127552 GCA_000007445.1 5M
D7 S. aureus GridION R10.4 e8.1 (FLO-MIN112) 450 4000 Dorado SUP v0.5.3 407,727 1,281M SRR21386013 GCF_000144955.2 2.8M
@@ D.2 Parameters
D.3 Versions
Supplementary Table S9 shows the version and the link to these corresponding versions of each tool and library we use in our experiments and in RawHash2, respectively.